# FlowMotion: Target-Predictive Flow Matching for Realistic Text-Driven Human Motion Generation

### v1

@@ 3.4 Sampling
$x_{i+1}=x_{i}-h\left(u_{t_{i}}(x_{i}|x_{1})\right),$ (6)
where $i\in\{0,1,\dots,M-1\}$, $M$ denotes the number of iterations, $h=1/M$ represents the step size, and the time instance is given by $t_{i}=i/M$. Observe that $t_{i}\in[0,1)$. The conditional vector field $u_{t_{i}}$, defined in Eq. (4), depends on $x_{1}$ and $x_{i}$.
@@
$\mathcal{G}(x_{t},t,c;\theta)=\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)+s\left(\mathcal{G}_{p}(x_{t},t,c;\theta)-\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)\right),$
@@
$x_{i+1}=x_{i}-h\Bigg{[}\frac{\mathcal{G}(x_{i},t_{i},c;\theta)-(1-\sigma_{\text{min}})x_{i}}{1-(1-\sigma_{\text{min}})t_{i}}\Bigg{]}.$
As previously mentioned, the objective of flow matching, and in our specific case CFM, is to determine the conditional vector field $u_{t}$. This is achieved from the model designed to estimate the target directly. By obtaining $\mathcal{G}$ and having the previous iterative value $x_{i}$, we can derive the desired vector field using Eq. (4) without the need to generate the vector field directly.
@@ 3.5 Implementation Details
During sampling, we set $\sigma_{min}$ to zero. We use Euler integration with 100 steps and a classifier-free guidance scale of 2.5, a standard value in similar works [20]. For optimization, we adopted the AdamW [61] optimizer, configured with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999, and a learning rate set to $1\times 10^{-4}$.
@@ 4 Experiments
All experiments were conducted on a Linux system equipped with an NVIDIA A40 GPU card. The training process typically took about 20 hours to complete, whereas generating samples required approximately 2 seconds per instance. This represents a significant speed advantage compared to diffusion models as MDM [20], which typically require about 30 seconds for sampling, and also surpasses other existing diffusion-based models in terms of sampling speed [19].
@@ 4.1 Datasets
    HumanML3D: This dataset, introduced by Guo et al. [30], combines the HumanAct12 [62] and AMASS [63] datasets. It consists of 14,616 motion sequences at 20 frames per second. The original motion sequences can exceed 10 seconds in duration, but following prior work [30], they are randomly truncated to a maximum of 10 seconds (196 frames) for consistency, with a minimum sequence length of 40 frames. The dataset is paired with 44,970 textual descriptions. HumanML3D is generally considered the primary benchmark for assessing the performance of different approaches in this task. The results of our experiments on this dataset are detailed in Table 1.
@@
Methods RP Top3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MMModality $\uparrow$ Jitter $\rightarrow$
Real motion $0.784^{\pm.003}$ $0.026^{\pm.003}$ $2.772^{\pm.013}$ $11.016^{\pm.095}$ $49.922$^{\pm$0.165$}$
T2M-GPT [21] $0.737^{\pm.004}$ $0.469^{\pm.010}$ $3.002^{\pm.013}$ $11.006^{\pm.111}$ $1.903^{\pm.070}$ $98.425$^{\pm$0.291$}$
MDM [20] $0.731^{\pm.004}$ $0.505^{\pm.027}$ $3.077^{\pm.017}$ $10.705^{\pm.098}$ $1.782^{\pm.152}$ $73.21^{\pm.5}$
MFM [39] $0.405^{\pm.004}$ $0.327^{\pm.016}$ $9.155^{\pm.028}$ $10.707^{\pm.070}$ $1.639^{\pm.137}$ $1099.988$^{\pm$1.183$}$
FlowMotion (ours) $0.404^{\pm.005}$ $0.396^{\pm.042}$ $9.206^{\pm.022}$ $10.989^{\pm.082}$ $1.756^{\pm.083}$ $52.40^{\pm.3}$
@@ 4.3 Metrics
The evaluation metrics employed are those proposed by Guo et al. [30], which have become standard practice in the human motion generation literature [21, 19, 55, 39, 20]. According to established evaluation protocols, such as those used by Guo et al. [30], each metric is computed over 20 trials to account for the inherent stochasticity in the sampling and evaluation process, and we report the average and standard deviation across these trials, as shown in Tables 12. The metrics used to evaluate 3D human motion generation are as follows:
Fréchet Inception Distance (FID)
Quantifies the similarity between the distribution of generated motions and the distribution of ground-truth motions from the test set, where both sets of motions are conditioned on the same textual descriptions from the test set.
@@
Measures the correspondence between the generated motion and the input text. We report Top-1, Top-2, and Top-3 R-Precision, representing the probability of retrieving the correct motion given the text within the top 1, 2, or 3 generated samples, respectively.
@@
Determines the average Euclidean distance between generated motions encoded in the same latent space when conditioned on the identical text.
@@ 4.4 Model Parameter Analysis
The scaled-up FlowMotion-Big variant improves the FID-Jitter balance marginally (Table 1); however, this comes at the cost of increased model complexity and computational overhead. Both variants retain the original Transformer 512-dimensional hidden representation in the encoder, consistent with prior work [64].
Figure 4: Comparison of motion trajectories generated by MDM [20], MFM [39], and the FlowMotion Model. MDM and MFM fail to depict a complete circular trajectory accurately. In contrast, the FlowMotion Model successfully generates a motion sequence in which the starting and ending points coincide. This result demonstrates the superior ability of the FlowMotion Model to interpret the motion instruction and generate a complete circular trajectory.
@@ 5 Results
Our model achieves optimal balance between FID and Jitter, demonstrating exceptional Jitter performance that approximates lifelike human motion. It outperforms all other methods in this regard, except for the diffusion model MDM, which exhibits a slightly better Jitter score but suffers from a higher FID and a significantly slower sampling speed (Tables 12). This balance suggests that our model is especially suitable for applications requiring both performance and high-quality motion generation.
On the HumanML3D dataset, our model achieves better FID than both the diffusion-based method MDM and the flow matching method MFM. While some models, such as MLD [36] and MotionLCM [60], achieve lower Jitter scores than ours, they achieve this with a substantial compromise in generalization capability, reflected in higher FID values (0.41 compared to our 0.27, Table 1). The lower Jitter scores, surpassing those of the ground truth, can potentially be attributed to factors such as differences in the generated sequence length or interpolation strategies, which may result in artificially smoother transitions. Conversely, models that exhibit better generalization than ours, such as the Vector Quantized Variational Autoencoder (VQ-VAE) based T2M-GPT and MotionGPT, tend to produce less smooth motion sequences, with Jitter values approximately double that of the ground truth (Tables 1-2).
Notably, our model maintains a low FID, indicative of strong generalization, and remarkably low Jitter scores on the HumanML3D dataset. If the goal is to generate motions as smooth as those in the ground truth data, our model achieves the second-best Jitter score, with MDM being the only model with a slightly lower Jitter (46 compared to the ground truth value of 47). However, MDM exhibits a much higher FID (0.52 compared to our 0.27), indicating a deficiency in generalization that our model successfully addresses (Fig. 4).
On the KIT dataset, our method achieves FID values generally around 0.4, coupled with the best Jitter score compared to all methods (Table 2), highlighting its ability to generate smooth, high-quality motion sequences while maintaining excellent generalization capability across different datasets. This superior balance is also visually apparent when comparing the generated motions, as shown in Fig. 5, which depicts a motion sequence from a side profile. As shown in Fig. 5, MFM produces motions with noticeable jerkiness. Our method, while having the same number of frames as MFM, exhibits smoother transitions throughout the motion, making it appear more natural. While matching MFM in the number of frames, our approach yields significantly more natural transitions, underscoring its effectiveness in generating high-fidelity motion.
@@
Further evidence of robustness of our model, when trained on KIT-ML [44], is presented in Figure 6, directly comparing motion generation against MFM [39]. A key advantage, enhanced motion stability, is visually highlighted: Figure 6a shows MFM generation with noticeable tremors and motion artifacts, including body part instability, even during intended stillness before a step. In contrast, captured at the same instant (Figure 6b), our FlowMotion model showcases stability, maintaining a coherent pose without involuntary movements or tremors. This visual comparison underscores improved stability and reduced artifact generation achieved by FlowMotion.
@@
We also analyze the trade-off between motion quality and sampling steps, as illustrated in Fig. 7, which shows that both FID and Jitter values exhibit a trend towards stabilization after approximately 50 sampling steps, approaching the error region of our previously reported results (horizontal lines in Fig. 7). A more complete set of metrics, including diversity and R-precision, is presented in Table 3 for sampling steps ranging from 1 to 100. Notably, all metrics remain relatively stable from around 50 steps onwards, consistent with our results in Table 1. This analysis, conducted on the HumanML3D test set, indicates that a sampling regime of around 50 steps is sufficient for generating high-quality motions with good generalization, striking a balance between computational cost and motion fidelity.
@@
Moreover, our model demonstrates a more precise interpretation of textual prompts when compared to MDM and MFM (Fig. 4). Notably, it successfully generates a complete circular trajectory with coinciding start and end points, a feat not achieved by MDM or MFM.
@@ 6 Conclusions
We introduce FlowMotion, a text-to-motion generation framework leveraging Conditional Flow Matching (CFM). Our key contribution is a novel training objective that directly predicts target motions, bypassing intermediate velocity estimation to reduce jitter and improve stability. Evaluations on the HumanML3D and KIT datasets demonstrate that FlowMotion achieves state-of-the-art fidelity (FID) and smoothness (jitter) while maintaining competitive efficiency against diffusion-based and flow matching baselines. The sampling process generates motion sequences in approximately 2 seconds—significantly faster than traditional diffusion models—enabling practical deployment in resource-constrained scenarios.
FlowMotion balances computational efficiency and motion quality without relying on oversized architectures. While some methods achieve marginally lower jitter, our framework prioritizes a holistic trade-off between perceptual realism, temporal consistency, and speed. Future work will explore enforcing physical plausibility (e.g., ground contact constraints to eliminate floating artifacts) and extending the framework to tasks like motion editing and full-body articulation. The proposed training objective also opens avenues for broader applications in flow-matching frameworks beyond human motion synthesis.
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) - Finance Code 001.

### v2

@@ 3.4 Sampling
$x_{t_{i+1}}=x_{t_{i}}+h\left(u_{t_{i}}(x_{t_{i}}|x_{1})\right),$ (6)
where $i\in\{0,1,\dots,M-1\}$, $M$ denotes the number of iterations, $h=1/M$ represents the step size, and the time instance is given by $t_{i}=i/M$. Observe that $t_{i}\in[0,1)$. The conditional vector field $u_{t_{i}}$, defined in Eq. (4), depends on $x_{1}$ and $x_{t_{i}}$.
@@
$\begin{split}\mathcal{G}(x_{t},t,c;\theta)&=\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)\\ &+s\left(\mathcal{G}_{p}(x_{t},t,c;\theta)-\mathcal{G}_{p}(x_{t},t,\emptyset;\theta)\right)\end{split}$
@@
$x_{t_{i+1}}=x_{t_{i}}+h\Bigg{[}\frac{\mathcal{G}(x_{t_{i}},t_{i},c;\theta)-(1-\sigma_{\text{min}})x_{t_{i}}}{1-(1-\sigma_{\text{min}})t_{i}}\Bigg{]}.$
As previously mentioned, the objective of flow matching, and in our specific case CFM, is to determine the conditional vector field $u_{t}$. This is achieved from the model designed to estimate the target directly. By obtaining $\mathcal{G}$ and having the previous iterative value $x_{t_{i}}$, we can derive the desired vector field using Eq. (4) without the need to generate the vector field directly.
@@ 3.5 Implementation Details
In our implementation, we set $\sigma_{min}$ to zero during both sampling and training. While theoretically, the standard deviation $\sigma_{min}$ of a Gaussian distribution must be strictly positive, empirically, setting $\sigma_{min}=0$ yielded improved performance in our experiments.
We use Euler integration with 100 steps and a classifier-free guidance scale of 2.5, a standard value in similar works [31]. For optimization, we adopted the AdamW [63] optimizer, configured with parameters $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999, and a learning rate set to $1\times 10^{-4}$.
@@ 4 Experiments
Experiments were conducted on a Linux system using NVIDIA A30 and RTX 3060 GPU cards. The A30 GPU was used for training, which took approximately 24 hours. The RTX 3060 GPU was used for evaluating metrics and comparing against state-of-the-art methods. Sample generation for our model, performed on the RTX 3060, required approximately 2 seconds per instance. In contrast, the MDM [31], which is based on diffusion, required approximately 30 seconds for sample generation on the same RTX 3060 GPU. This speed constitutes a significant improvement and also surpasses other diffusion-based models in sampling speed [32].
@@ 4.1 Datasets
    HumanML3D: This dataset, introduced by Guo et al. [26], combines the HumanAct12 [64] and AMASS [65] datasets. It consists of 14,616 motion sequences at 20 frames per second. The original motion sequences can exceed 10 seconds in duration, but following prior work [26], they are randomly truncated to a maximum of 10 seconds (196 frames) for consistency, with a minimum sequence length of 40 frames. The dataset is paired with 44,970 textual descriptions. The results of our experiments on this dataset are detailed in Table 1.
@@
Methods #params $\downarrow$ FID $\downarrow$ Diversity $\uparrow$ Jitter $\rightarrow$ FID-J $\downarrow$ MM-Dist $\downarrow$ MMModality $\uparrow$ RP Top3 $\uparrow$
Real motion - $0.026^{\pm.003}$ $11.016^{\pm.095}$ $49.922$^{\pm$0.165$}$ $0.00$ $2.772^{\pm.013}$ $0.784^{\pm.003}$
T2M-GPT [24] 228.4M $0.469^{\pm.010}$ $11.006^{\pm.111}$ $98.425$^{\pm$0.291$}$ $8.729$ $3.002^{\pm.013}$ $1.903^{\pm.070}$ $0.737^{\pm.004}$
MDM [31] 17.9M $0.505^{\pm.027}$ $10.705^{\pm.098}$ $73.21^{\pm.5}$ $10.196$ $3.077^{\pm.017}$ $1.782^{\pm.152}$ $0.731^{\pm.004}$
MFM [39] 17.9M $0.327^{\pm.016}$ $10.707^{\pm.070}$ $1099.988$^{\pm$1.183$}$ $53.801$ $9.155^{\pm.028}$ $1.639^{\pm.137}$ $0.405^{\pm.004}$
FlowMotion 17.9M $0.396^{\pm.042}$ $10.989^{\pm.082}$ $52.40^{\pm.3}$ $8.362$ $9.206^{\pm.022}$ $1.756^{\pm.083}$ $0.404^{\pm.005}$
@@ 4.3 Metrics
We employ evaluation metrics suggested by Guo et al. [26], a standard practice in the human motion generation literature [24, 32, 58, 39, 31]. These metrics include Fréchet Inception Distance (FID), R-Precision (RP Top-k), Multimodal Distance (MM-Dist), Diversity, and Multimodality (MModality). Following established evaluation protocols, each metric, except for MModality, is computed over 20 trials to account for the inherent stochasticity in the sampling and evaluation process. MModality is calculated after generating 30 samples for a same set of randomly chosen batches. We report the average and standard deviation across these trials, as shown in Tables 12. In addition to this established suite, we introduce Jitter, which quantifies motion smoothness, and Mahalanobis FID-J, which measures the balance between fidelity and motion naturalness. As with the other metrics, Jitter is computed over 20 trials to account for stochasticity.
@@
Measures the correspondence between the generated motion and the input text. For RP Top-k, batches of 32 (text embedding, motion embedding) pairs from the test set are processed. The distance between each text embedding and all motion embeddings in the batch is calculated. We count how often the correct motion embedding appears among the 1, 2, or 3 nearest neighbors (Top-1, Top-2, Top-3). This is repeated for all batches, and the final percentage is calculated by dividing by the total number of samples.
@@
Determines the average Euclidean distance between generated motions encoded in the same latent space when conditioned on the identical text. MModality is calculated after generating 30 samples for the same set of randomly chosen batches.
Params
reports the total number of parameters in each model’s architecture. It reflects the overall size and complexity of the network used during training.
@@
Mahalanobis FID-J
We introduce the Mahalanobis FID-J metric to quantify the balance between FID and motion jitter. It measures the Mahalanobis distance, defined as:
$D_{M}(x)=\sqrt{(x-\mu)^{T}S^{-1}(x-\mu)},$ (7)
of a method’s FID and jitter vector ($x$) from the ground truth FID and jitter vector ($\mu$) in the FID-Jitter space. In this context, $S$ is the covariance matrix. The ground truth values represent the ideal performance point, and minimizing the Mahalanobis distance implies achieving an optimal balance between these two metrics. Given that FID and Jitter have different scales (Fig. 1), the Mahalanobis distance, a statistical measure that accounts for the covariance between the metrics, is more appropriate than the Euclidean distance. This allows us to effectively evaluate how well a method jointly minimizes both FID and jitter.
To ensure a fair comparison of methods, outlier detection was applied to the FID and Jitter values solely for the purpose of computing the covariance matrix ($S$ in Equation 7), particularly due to the presence of high Jitter values for the MFM method (576.68 and 1099.99) in Table 1 (HumanML3D) and Table 2 (KIT), respectively.
  • 1.
    For HumanML3D (Table 1), outlier detection was performed using the Mahalanobis distance, identifying the TM2T method (FID=1.507, Jitter=87.21) and the MFM method (FID=0.446, Jitter=576.68) as the outlier.
  • 2.
    For the KIT dataset (Table 2), given the limited number of evaluated models (four), we employed a robust z-score based on the Median Absolute Deviation (MAD), applied specifically to the jitter values to identify potential outliers. This identified the MFM method with a Jitter value of 1099.99 as an outlier. Values with $|z|>3$ were considered outliers.
For both HumanML3D (Table 1) and KIT (Table 2), after computing the covariance matrix (after outlier removal if applicable), this matrix was then used to calculate the Mahalanobis distance between all methods and the ground truth reference, using Equation 7, allowing for a more robust and reliable comparison across methods.
@@ 4.4 Model Parameter Analysis
The scaled-up FlowMotion-Big variant improves the FID-Jitter balance marginally (Table 1); however, this comes at the cost of increased the number of parameter of the model. Both variants retain the original Transformer 512-dimensional hidden representation in the encoder, consistent with prior work [66].
Figure 4: Comparison of trajectory generation performance between MDM [31], MFM [39], and the FlowMotion Model for circular motion instructions. MDM and MFM exhibit significant errors in trajectory closure. The FlowMotion Model, conversely, generates a trajectory with accurate start and end point coincidence, indicative of superior motion instruction understanding. The FlowMotion Model’s improved FID score confirms its enhanced generation fidelity to the intended circular motion.
@@ 5 Results
Our model demonstrates a superior trade-off between fidelity and motion smoothness, achieving notably low Jitter values while maintaining high-quality motion generation. On the KIT dataset (Table 2), it achieves the best Jitter reduction with a score of 52.4, close to the ground truth value of 49.9. On HumanML3D (Table 1), it achieves a Jitter score of 39.52, outperforming MDM (46.3) and approaching the ground truth (47.26). While MDM shows a slightly better Jitter on HumanML3D, it suffers from significantly higher FID and a substantially slower sampling speed (30 seconds vs. our 2 seconds).
To better capture the trade-off between fidelity and naturalness, we introduce the FIDJ metric, demonstrating that we achieve an optimal balance on both datasets. This metric measures the Mahalanobis distance from each method’s (FID, Jitter) pair to the ground truth point: (0.001, 47.26) for HumanML3D and (0.026, 49.92) for KIT. In this formulation, the ground truth naturally has a FIDJ of 0, as it defines the reference point. Our model achieves FIDJ scores of 2.4 on HumanML3D and 8.3 on KIT, indicating that it remains closest to the optimal point in the FID–Jitter space. These results highlight the model’s effectiveness for applications requiring both realistic and high-fidelity human motion generation.
On the HumanML3D dataset, our model—built upon the same underlying architecture as MDM and MFM—achieves a lower FID score than both (Fig. 4). In contrast, models using Vector Quantized Variational Autoencoders (VQ-VAE), such as T2M-GPT and MotionGPT, exhibit higher fidelity (FID) but tend to produce less smooth motion sequences, with Jitter values approximately double that of the ground truth (Tables 1-2). The superior balance between motion quality and smoothness is also visually apparent when comparing the generated motions, as shown in Fig. 5, which depicts a motion sequence from a side profile. As shown in Fig. 5, MFM produces motions with noticeable jerkiness. Our method, while having the same number of frames as MFM, exhibits smoother transitions throughout the motion, making it appear more natural, underscoring its effectiveness in generating high-fidelity motion.
@@
On the KIT dataset, our method achieves a low FID-J (8.3) that outperforms all other methods in Table 2, highlighting its ability to generate smooth, high-quality motion sequences while maintaining excellent fidelity. Further evidence of robustness of our model, when trained on KIT-ML [42], is presented in Figure 6, directly comparing motion generation against MFM [39]. A key advantage, enhanced motion stability, is visually highlighted: Figure 6a shows MFM generation with noticeable tremors and motion artifacts, including body part instability, even during intended stillness before a step. In contrast, captured at the same instant (Figure 6b), our FlowMotion model showcases stability, maintaining a coherent pose without involuntary movements or tremors. This visual comparison underscores improved stability and reduced artifact generation achieved by FlowMotion.
@@
5.1 Ablation Study
Regarding the ablation study, it is important to note that the original work of MFM achieved smoother results, albeit with a larger batch size (256). Due to hardware limitations, we were constrained to use a batch size of 128. Nevertheless, MFM represents the process of using flow matching training to determine the vector field; hence, it serves as our ablation study. In our case, we perform the training by directly determining the value without noise $x_{1}$. The results demonstrate the advantages of our direct training approach over the conventional flow-matching technique, even under similar training conditions.
@@ 5.1 Ablation Study
Our model also performs strongly on Diversity, achieving scores among the top two on both the HumanML3D (Table 1) and KIT (Table 2) datasets, suggesting that it explores the motion space in a rich and varied manner. Conversely, it shows lower performance on the RP Top3, MM-Dist, and MM-Modality metrics, indicating a less strict adherence to the text condition. This can be attributed to the model’s focus on preserving realism and diversity rather than replicating canonical or literal motions. We consider this trade-off consistent with the model’s objectives, which aim to offer more fluid, natural, and diverse generations, even if they are less deterministic.
We also analyze the trade-off between motion quality and sampling steps, as illustrated in Fig. 7, which shows that both FID and Jitter values exhibit a trend toward stabilization after approximately 50 sampling steps, approaching the error region of our previously reported results (horizontal lines in Fig. 7). A more complete set of metrics, including diversity and R-precision, is presented in Table 3 for sampling steps ranging from 1 to 100. Notably, all metrics remain relatively stable from around 50 steps onward, consistent with our results in Table 1. This analysis, conducted on the HumanML3D test set, indicates that a sampling regime of around 50 steps is sufficient for generating high-quality motions with good fidelity, striking a balance between computational cost and motion fidelity.
@@ 6 Conclusions
We introduced FlowMotion, a text-to-motion generation framework leveraging Conditional Flow Matching (CFM). To achieve highly diverse and perceptually consistent 3D character animations with natural motion and low computational costs, FlowMotion incorporates a training objective within CFM that more accurately predicts target motion. We apply this target prediction approach to 3D human motion generation to reduce the inherent jitter associated with CFM while enhancing stability, realism, and computational efficiency. Evaluations on the HumanML3D and KIT datasets demonstrated that FlowMotion achieves a superior balance between fidelity (FID) and smoothness (jitter) while maintaining competitive efficiency against diffusion-based and flow matching baselines. Our framework excels in achieving low jitter values, particularly in the KIT dataset, while also maintaining very competitive FID scores, leading to an excellent Mahalanobis FID-J metric. The sampling process generates motion sequences with the fast generation speeds characteristic of flow matching, a significant improvement over traditional diffusion models, enabling practical deployment in resource-constrained scenarios.
FlowMotion balances computational efficiency and motion quality without relying on oversized architectures. While some methods achieve marginally lower Jitter, our framework prioritizes a holistic trade-off between perceptual realism, temporal consistency, and generation speed, as evidenced by our strong Diversity scores. While our model excels in generating diverse and natural human motions, future research will investigate methods to enhance text-conditional fidelity, specifically targeting improvements in RP Top3, MM-Dist, and MM-Modality metrics. We also plan to extend the framework to applications such as interactive motion editing and full-body articulation, leveraging the model’s inherent ability to produce smooth and varied movements.
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.